# Channel Choice and Customer Value

### v1

@@ E.2 Appendix to the Other Metrics
Table A25 and A26 presents DID results for the total and offline metrics under the covid analysis.
@@
Table A27 and A28 presents the DID estimates of other total metrics and offline metrics of interest under the Black Friday analysis.
@@
Table A29 and A30 presents the DID estimates of the other total metrics and offline metrics of interest under the loyalty program analysis.
@@
E.3 Appendix to Forward buying
@@
We first compare the spend of organic and loyalty program adopters during their adoption month relative to the pre-adoption reference period (i.e., spend in May 2023), and Table A32 shows that loyalty program adopters tend to spend 177.7 MCU more than organic adopters during the adoption month. Next, we notice that both loyalty program and organic adopters can activate the monthly loyalty offer during the post-adoption period. To test the post-adoption differences in purchase behavior, we estimate the event study regression in Equation (5) with control variables for the customers’ activations. Table A33 shows that loyalty program adopters spend significantly less than organic adopters in most of the post-adoption periods.
@@
E.4 Appendix to Managerial Implications
Table A34 presents an additional DID result for profit under Black Friday analysis, which are inputs for the discussion of managerial implications. Instead of estimating Equation (1) and (2), with a single indicator $\mathbbm{I}\{Adopter_{i}\}$, we now include two separate dummies – $\mathbbm{I}\{Event\_Adopter_{i}\}$ and $\mathbbm{I}\{Organic\_Adopter_{i}\}$ – each interacted with the post-adoption indicator. This allows us to estimate the ATT of online adoption on profit for organic adopters and event adopters, relative to offline-only customers.
Table A34: DID Results of Profit – Black Friday Adopters, Organic Adopters, and Offline-only Customers
DV: Profit
Model: (1) (2)
@@ E.4 Appendix to Managerial Implications
Constant 96.13 (0.187)
@@
Organic_Adopter 58.55 (3.523)
[0.000]
Post -22.83 (0.152)
[0.000]
Organic_Adopter * Post 26.14 (3.166) 23.57 (3.160)
@@
Observations 9,505,999 9,505,999
R2 0.00451 0.34750
Within R2 $2.88\times 10^{-5}$
while p‑values are shown in square brackets.
Appendix F Appendix to Validity and Robustness Checks
F.1 Appendix to Parallel Trend Tests
F.1.1 Parallel Trend Test for Covid Analysis - Total Spend
@@
Then, Figure 6(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for the total spend in Web Appendix §F.2. The incremental spend due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 6(c) presents the event study estimated coefficients between organic adopters and covid adopters. As we can see, the estimated coefficients are not significantly different from zero, i.e., the parallel pre-trend assumption seems valid.  [^36] For the parallel trend assumption to hold, it is permissible for a small number of pre-treatment coefficients to be statistically significant (provided the majority of these coefficients are statistically insignificant), see Gopalan, Rosinger, and Ahn (2020) and Chapter 5 of Angrist and Pischke (2008). This suggests that the DID estimation of treatment effects was not falsely inflated by any trends before the treatment. As a further robustness check, we also provide bounded ATT estimates for the total spend result between organic adopters and covid adopters in Web Appendix §F.2.
F.1.2 Parallel Trend Tests for Covid Analysis - Profitability Analysis
Figure A7 presents the parallel trend plot for the profit. Figure 7(a) visualizes the trends. Figure 7(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for profit in Web Appendix §F.2. The incremental profit due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 7(c) confirms that the parallel trend assumption for total profit between covid adopters and organic adopters is satisfied since almost all pre-treatment coefficients are insignificant and not statistically different from 0. As a further robustness check, we also provide bounded ATT estimates for the profit result between organic adopters and covid adopters in Web Appendix §F.2.
@@ F.1.2 Parallel Trend Tests for Covid Analysis - Profitability Analysis
F.1.3 Additional Parallel Trend Plots for Covid Analysis – Other Metrics
In this section, we validate the parallel trend assumption for DID results in other metrics between organic adopters and covid adopters, presented in Table A25 and A26. Because both organic and covid adopters only purchase offline during the pre-adoption period, the pre-adoption values for total and offline metrics are exactly the same. So here we test for total metrics and same conclusions apply to offline metrics as well. In Figure A8, we present the parallel trend plots for the metrics shown in Table A25. Figure A9 shows the event study estimates test for pre-trend. We can see some significant pre-treatment coefficients for these metrics, but the magnitude of these coefficients are small.
@@ F.1.3 Additional Parallel Trend Plots for Covid Analysis – Other Metrics
F.1.4 Parallel Trend for Black Friday Analysis
First, we provide a pre-trend test for two main metrics (i.e. spend and profit) between adopters (including both Black Friday adopters and Organic adopters) relative to offline-only customers. In total, we have 10 pre-treatment periods (i.e., between January 2019 and October 2019), labeled as -10 to -1, and 11 post-adoption periods (i.e., between December 2019 and October 2020), labeled as 1 to 11. The base period is Octorber 2019, and the event time 0 refers to the adoption month (i.e., November 2019). We exhibit a minor violation during the early stage of the pre-adoption period, but most pre-adoption coefficients are significant. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §F.2 and confirm that the incremental spend and profit are still positive and significant in the presence of violation of parallel trend.
@@ F.1.4 Parallel Trend for Black Friday Analysis
F.1.5 Parallel Trend for Loyalty Program Analysis
First, we provide a pre-trend test for two main metrics (i.e. spend and profit) between adopters (including both Loyalty Program adopters and Organic adopters) relative to offline-only customers (see Figure A13). We exhibit a violation of parallel trend assumption. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §F.2 and confirm that the incremental spend is still positive and significant in the presence of violation of parallel trend.
@@ F.1.5 Parallel Trend for Loyalty Program Analysis
Next, we provide a pre-trend test between Loyalty Program adopters and Organic adopters. Figure A14 visualizes the monthly average metric values for loyalty program adopters, organic adopters and offline-only customers. Figure A15 present the event study regression estimates for the parallel pre-trend of the loyalty program online activation study.
@@
F.2 Honest DID
In F.1, we see that when comparing adopters and offline-only customers, there are some violations of parallel trend assumption. There are also some concerns about the parallel trend assumption between covid adopters and organic adopters. To ensure that estimated ATTs of online adoption are robust to the violations in the parallel trends assumption, we perform the Honest Difference-in-Differences (Honest DID) estimation to quantify the extent to which violations in parallel trends can nullify the treatment effect (Rambachan and Roth 2023).
@@ F.2.3 Loyalty Program
F.3 Appendix to DID Models with Propensity Score Matching and Inverse Propensity Score Weighting
In this section, we provide a set of robustness check under PSM and IPTW. To conduct this analysis, we label the event adopters as 1 and organic adopters as 0 and model the propensity score of customer $i$ being a COVID adopter. The propensity score model is specified as follows:
@@ F.3 Appendix to DID Models with Propensity Score Matching and Inverse Propensity Score Weighting
[…]
Next, we use these propensity scores to augment our DID anaysis in two ways.
@@
    First, we perform a customer-level propensity score matching and then re-run the DID analysis. Matching based on propensity score is much more efficient than matching based on a set of covariates (Rosenbaum and Rubin 1985). Based on the predicted propensity score of each customer $i$, we use MatchIt package in R (Stuart et al. 2011) to perform nearest-neighbor matching for each event adopter. We then run the DID analysis on the newly matched sample (see Table A41, A42 and A43 in Web Appendix F.3.2). We find that the results from this exercise are consistent with the main findings.
@@
    Next, we use a slightly different approach to control for any potential selection issues. Specifically, we use an inverse propensity of treatment weight-adjusted (IPTW) DID regression. Here, the sample is the same as the one from the main analysis, but each observation is inversely weighted by its propensity score. We then perform a two-sample t-test on all the pre-treatment variables based on the weighted sample and confirm that the two samples are statistically insignificant on all the variables (see Table A40 for the weighted sample of organic adopters and event adopters, respectively in the Web Appendix F.3.3). The DID results of event adopters and organic adopters from this analysis are shown in Table A44, A45 and A46 for covid, Black Friday, and Loyalty Program analysis, respectively in Web Appendix F.3.3. Again, we find that the results and conclusions are quite similar to those in the main section.
F.3.1 Propensity Score Model Estimation
@@
Table A39 presents the propensity score model estimates for covid, black friday, and loyalty program adopters versus organic adopters. Figure A20 presents the distribution of propensity scores of event adopters and organic adopters under three studies. After performing nearest-neighbor matching, A21 shows that the distribution of propensity scores between event adopters and organic adopters become more overlapped.
@@
F.3.2 DID Results with PSM
Table A41, A42 and A43 present the DID estimates of Equation (3) of key purchase metrics based on the matched sample under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
@@ F.3.2 DID Results with PSM
F.3.3 DID Results with IPTW
Table A44, A45 and A46 present the DID estimates of Equation (3) of key purchase metrics with inverse propensity score weighting under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
@@ F.3.3 DID Results with IPTW
F.3.4 Doubly Robust DID for Covid Analysis
@@
F.4 Appendix to Alternate definitions for the organic adopter and covid adopter cohorts
F.4.1 Appendix to Covid Analysis - Re-defining the Organic Adopter Cohort
In this section, we provide a robustness check by redefining organic adopters as those customers who were offline-only before 2019-10-31 and made their first online purchase between 2019-11-01 to 2020-02-29. The definition of covid adopters remains the same. Under this new definition, we continue to have 33,246 covid adopters and a larger number of 25,958 organic adopters. As before, we exclude the periods used for defining these two cohorts (2019-11-01 to 2020-04-30) and re-run the same set of DID models as in the main analysis. Table A48 presents the DID results between covid adopters and organic adopters. Table A49 and A50 presents DID results for other total and offline metrics respectively. We find that all the results are consistent with the main analysis.

### v2

@@ D.2 Appendix to the Other Metrics
Tables A19 and A20 present DID results for the total and offline metrics under the covid analysis.
@@
Tables A21 and A22 present the DID estimates of other total metrics and offline metrics of interest under the Black Friday analysis.
@@
Tables A23 and A24 present the DID estimates of the other total metrics and offline metrics of interest under the loyalty program analysis.
@@
D.3 Appendix to Forward buying
@@
We first compare the spend of organic and Loyalty Program adopters during their adoption month relative to the pre-adoption reference period (i.e., spend in May 2023), and Table A26 shows that Loyalty Program adopters tend to spend 177.7 MCU more than organic adopters during the adoption month. Next, we notice that both Loyalty Program adopters and organic adopters can activate the monthly loyalty offer during the post-adoption period. To test the post-adoption differences in purchase behavior, we estimate the event study regression in Equation (5) with control variables for the customers’ activations. Table A27 shows that Loyalty Program adopters spend significantly less than organic adopters in most of the post-adoption periods.
@@
D.4 Appendix to Managerial Implications
Table A28 presents an additional DID result for profit, which are inputs for the discussion of managerial implications. Instead of estimating Equation (1) and (2), with a single indicator $\mathbbm{I}\{Adopter_{i}\}$, we now include two separate dummies – $\mathbbm{I}\{Event\_Adopter_{i}\}$ and $\mathbbm{I}\{Organic\_Adopter_{i}\}$ – each interacted with the post-adoption indicator. This allows us to estimate the ATT of online adoption on profit for organic adopters and event adopters, relative to offline-only customers.
Table A28: DID Results (All Three Groups) – Profit
Study: Covid Black Friday Loyalty Program
Dependent Var.: Profit
Model: (1) (2) (3) (4) (5) (6)
@@ D.4 Appendix to Managerial Implications
Constant 92.57 (0.177) 96.13 (0.187) 125.8 (0.288)
[0.000] [0.000] [0.000]
Organic_Adopter 53.18 (1.707) 58.55 (3.523) 38.65 (1.652)
[0.000] [0.000] [0.000]
Covid_Adopter 64.14 (1.085)
@@
Post -10.56 (0.174) -22.83 (0.152) -16.81 (0.238)
[0.000] [0.000] [0.000]
Organic_Adopter * Post 30.83 (1.796) 30.28 (1.903) 26.14 (3.166) 23.57 (3.160) -4.244 (1.803) 18.55 (1.509)
[0.000] [0.000] [0.000] [0.000] [0.019] [0.000]
Covid_Adopter * Post 38.08 (1.147) 39.78 (1.272)
@@
Loyalty_Adopter 72.78 (6.210)
[0.000]
Loyalty_Adopter * Post -91.07 (7.814) 2.986 (6.798)
[0.000] [0.292]
Loyalty program controls Yes Yes
@@
R2 0.01237 0.31903 0.00451 0.34750 0.00660 0.44546
Within R2 0.00047 $2.88\times 10^{-5}$ 0.00194
Additional Analysis for ROI/breakeven analyses for promotion:
[…]
When the cost per customer is 50 MCU, if managers naively used the incremental profit from organic adopters, they would compute a breakeven time of 2.70 months (i.e., 50/18.55), whereas the correct breakeven period using Loyalty Program adopters’ incremental profit is 16.74 months (i.e., 50/2.986) – more than six times longer. Therefore, managers should appropriately calibrate their expectations of increased consumer spending when running promotions that encourage the adoption of online shopping.
When the cost per customer is 150 MCU, if managers naively used the incremental profit from organic adopters, they would compute a breakeven time of 8.09 months (i.e., 150/18.55), whereas the correct breakeven period using Loyalty Program adopters’ incremental profit is 50.23 months (i.e., 150/2.986) – more than six times longer. Therefore, managers should appropriately calibrate their expectations for increased consumer spending when running promotions that encourage online shopping adoption.
Appendix E Appendix to Validity and Robustness Checks
E.1 Appendix to Effect Size Measurement
Figure A4 presents the bootstrapped distribution of true ATT for covid adopters relative to organic adopters. Figure A5 presents the bootstrapped distribution of true ATT for Black Friday adopters relative to organic adopters.
Figure A4: True ATT – Covid Adopters vs. Organic Adopters
(a) Spend
(b) Profit
Figure A5: True ATT – Black Friday Adopters vs. Organic Adopters
(a) Spend
(b) Profit
E.2 Appendix to Parallel Trend Tests
E.2.1 Parallel Trend Test for Covid Analysis - Total Spend
@@
Then, Figure 6(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for the total spend in Web Appendix §E.3. The incremental spend due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 6(c) presents the event study estimated coefficients between organic adopters and covid adopters. As we can see, the estimated coefficients are not significantly different from zero, i.e., the parallel pre-trend assumption seems valid.  [^35] For the parallel trend assumption to hold, it is permissible for a small number of pre-treatment coefficients to be statistically significant (provided the majority of these coefficients are statistically insignificant), see Gopalan, Rosinger, and Ahn (2020) and Chapter 5 of Angrist and Pischke (2008). This suggests that the DID estimation of treatment effects was not falsely inflated by any trends before the treatment. As a further robustness check, we also provide bounded ATT estimates for the total spend result between organic adopters and covid adopters in Web Appendix §E.3.
E.2.2 Parallel Trend Tests for Covid Analysis - Profitability Analysis
Figure A7 presents the parallel trend plot for the profit. Figure 7(a) visualizes the trends. Figure 7(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for profit in Web Appendix §E.3. The incremental profit due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 7(c) confirms that the parallel trend assumption for total profit between covid adopters and organic adopters is satisfied since almost all pre-treatment coefficients are insignificant and not statistically different from 0. As a further robustness check, we also provide bounded ATT estimates for the profit result between organic adopters and covid adopters in Web Appendix §E.3.
@@ E.2.2 Parallel Trend Tests for Covid Analysis - Profitability Analysis
E.2.3 Additional Parallel Trend Plots for Covid Analysis – Other Metrics
In this section, we validate the parallel trend assumption for DID results in other metrics between organic adopters and covid adopters, presented in Table A19 and A20. Because both organic and covid adopters only purchase offline during the pre-adoption period, the pre-adoption values for total and offline metrics are exactly the same. So here we test for total metrics and same conclusions apply to offline metrics as well. In Figure A8, we present the parallel trend plots for the metrics shown in Table A19. Figure A9 shows the event study estimates test for pre-trend. We can see some significant pre-treatment coefficients for these metrics, but the magnitude of these coefficients are small.
@@ E.2.3 Additional Parallel Trend Plots for Covid Analysis – Other Metrics
E.2.4 Parallel Trend for Black Friday Analysis
First, we provide a pre-trend test for two main metrics (i.e., spend and profit) between adopters (including both Black Friday adopters and organic adopters) relative to offline-only customers. In total, we have 10 pre-treatment periods (i.e., between January 2019 and October 2019), labeled as -10 to -1, and 11 post-adoption periods (i.e., between December 2019 and October 2020), labeled as 1 to 11. The base period is October 2019, and the event time 0 refers to the adoption month (i.e., November 2019). We exhibit a minor violation during the early stage of the pre-adoption period, but most pre-adoption coefficients are significant. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §E.3 and confirm that the incremental spend and profit are still positive and significant in the presence of violation of parallel trend.
@@ E.2.4 Parallel Trend for Black Friday Analysis
E.2.5 Parallel Trend for Loyalty Program Analysis
First, we provide a pre-trend test for two main metrics (i.e., spend and profit) between adopters (including both Loyalty Program adopters and Organic adopters) relative to offline-only customers (see Figure A13). We exhibit a violation of the parallel trend assumption. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §E.3 and confirm that the incremental spend is still positive and significant in the presence of violation of parallel trend.
@@ E.2.5 Parallel Trend for Loyalty Program Analysis
Next, we provide a pre-trend test between Loyalty Program adopters and organic adopters. Figure A14 visualizes the monthly average metric values for Loyalty Program adopters, organic adopters, and offline-only customers. Figure A15 presents the event study regression estimates for the parallel pre-trend of the loyalty program online activation study.
@@
E.3 Honest DID
In E.2, we see that when comparing adopters and offline-only customers, there are some violations of the parallel trend assumption. There are also some concerns about the parallel trend assumption between covid adopters and organic adopters. To ensure that estimated ATTs of online adoption are robust to the violations in the parallel trends assumption, we perform the Honest Difference-in-Differences (Honest DID) estimation to quantify the extent to which violations in parallel trends can nullify the treatment effect (Rambachan and Roth 2023).
@@ E.3.3 Loyalty Program
E.4 Appendix to DID Models with Propensity Score Matching and Inverse Propensity Score Weighting
In this section, we provide a set of robustness checks under PSM and IPTW. To conduct this analysis, we label the event adopters as 1 and organic adopters as 0 and model the propensity score of customer $i$ being an event adopter. The propensity score model is specified as follows:
@@ E.4 Appendix to DID Models with Propensity Score Matching and Inverse Propensity Score Weighting
where $f(\cdot)$ is the logit link function where $f(X_{i}^{T}\theta)=\frac{1}{1+\exp(-X_{i}^{T}\theta)}$. $X_{i}\in\mathcal{R}^{M\times 1}$ is a column vector with $M$ covariates of customer $i$. These covariates include $\textit{Gender}_{i},\textit{Age}_{i},\textit{Tenure(Days)}_{i}$, $\textit{AverageSpendPerMonth}_{i}$, etc. $\theta\in\mathcal{R}^{M\times 1}$ is a column vector of coefficients. Table A33 in Web Appendix E.4.1 presents the propensity score model estimation for covid, Black Friday, and loyalty program analysis, respectively. Figure A20 presents the distribution of propensity scores before matching, and we observe some differences between event adopters and organic adopters. After matching, Figure A21 suggests that event adopters and organic adopters become more overlapped.
Next, we use these propensity scores to augment our DID analysis in two ways.
@@
    First, we perform a customer-level propensity score matching and then re-run the DID analysis. Matching based on propensity score is much more efficient than matching based on a set of covariates (Rosenbaum and Rubin 1985). Based on the predicted propensity score of each customer $i$, we use MatchIt package in R (Stuart et al. 2011) to perform nearest-neighbor matching for each event adopter. We then run the DID analysis on the newly matched sample (see Table A35, A36, and A37 in Web Appendix E.4.2). We find that the results from this exercise are consistent with the main findings.
@@
    Next, we use a slightly different approach to control for any potential selection issues. Specifically, we use an inverse propensity of treatment weight-adjusted (IPTW) DID regression. Here, the sample is the same as the one from the main analysis, but each observation is inversely weighted by its propensity score. We then perform a two-sample t-test on all the pre-treatment variables based on the weighted sample and confirm that the two samples are statistically insignificant on all the variables (see Table A34 for the weighted sample of organic adopters and event adopters, respectively in the Web Appendix E.4.3). The DID results of event adopters and organic adopters from this analysis are shown in Table A38, A39 and A40 for covid, Black Friday, and Loyalty Program analysis, respectively in Web Appendix E.4.3. Again, we find that the results and conclusions are quite similar to those in the main section.
E.4.1 Propensity Score Model Estimation
@@
Table A33 presents the propensity score model estimates for covid adopters, Black Friday adopters, and Loyalty Program adopters versus organic adopters. Figure A20 presents the distribution of propensity scores of event adopters and organic adopters under three studies. After performing nearest-neighbor matching, A21 shows that the distribution of propensity scores between event adopters and organic adopters becomes more overlapped.
@@
E.4.2 DID Results with PSM
Table A35, A36 and A37 present the DID estimates of Equation (3) of key purchase metrics based on the matched sample under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
@@ E.4.2 DID Results with PSM
E.4.3 DID Results with IPTW
Table A38, A39 and A40 present the DID estimates of Equation (3) of key purchase metrics with inverse propensity score weighting under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
@@ E.4.3 DID Results with IPTW
E.4.4 Doubly Robust DID for Covid Analysis
@@
E.5 Appendix to Alternate definitions for the organic adopter and covid adopter cohorts
E.5.1 Appendix to Covid Analysis - Re-defining the Organic Adopter Cohort
In this section, we provide a robustness check by redefining organic adopters as those customers who were offline-only before 2019-10-31 and made their first online purchase between 2019-11-01 to 2020-02-29. The definition of covid adopters remains the same. Under this new definition, we continue to have 33,246 covid adopters and a larger number of 25,958 organic adopters. As before, we exclude the periods used for defining these two cohorts (2019-11-01 to 2020-04-30) and re-run the same set of DID models as in the main analysis. Table A42 presents the DID results between covid adopters and organic adopters. Table A43 and A44 present DID results for other total and offline metrics, respectively. We find that all the results are consistent with the main analysis.